# NeRF-CA: Dynamic Reconstruction of X-ray Coronary Angiography with Extremely Sparse-views

### v1

@@ VI Discussion and Conclusion
We propose NeRF-CA, the first step toward a 4D Neural Radiance Field reconstruction method for X-ray coronary angiography in an extremely sparse-view setting. We propose a method that uniquely integrates scene decomposition with tailored regularization techniques to accurately decouple the CA scene in a static background and dynamic coronary artery structure. We demonstrate the effectiveness of our method both in the dynamic and sparse-view reconstruction settings. We introduce techniques such as static vs. dynamic factorization and dynamic entropy regularizers, weighted pixel sampling, and windowed positional encoding to ensure the separation of foreground and background. We demonstrate the effectiveness of windowed positional encoding and dynamic occlusion regularization to accurately reconstruct in the extremely sparse-view setting. We have compared NeRF-CA to existing NeRF-based techniques for sparse-view 3D reconstruction, significantly outperforming them for CA blood vessel reconstruction. Our ablation studies have also demonstrated the effectiveness of the individual components proposed in our method. Notably, we achieve useful 4D reconstructions with four angiogram sequences, aligning well with clinical workflows and showing potential for clinical adaptation.
@@
Despite these advantages, we also highlight a few limitations of NeRF-CA. Although showing its potential, NeRF-CA does not fully address all challenges in a clinical setting, such as motion or geometry inaccuracy in CA data. Motion-related challenges include respiratory motion, contrast inhomogeneity, and synchronization across viewpoints. Specifically, our method assumes synchronous motion with the ECG signal across viewpoints, which may no longer hold when incorporating these additional challenges. Therefore, we rely on synthetic CA data to perform quantitative evaluation. Future work should consider the various CA-based works proposed to deal with these challenges [3], although modifications may be needed to incorporate them into NeRF-CA. Additionally, geometry inaccuracies from C-arm calibration in clinical settings, unlike the perfect camera geometry assumed in our work, could impact performance and require further investigation. Finally, improving the running time of our method from hours to minutes is crucial for integration into clinical workflows. Future work could explore techniques like volumetric representations [37] or 3D Gaussian Splatting [51], which have shown promise in accelerating NeRF-based medical reconstructions.
@@
In this work, we proposed NeRF-CA, the first step toward a 4D reconstruction method for CA data in the extremely sparse-view setting without user interaction. We demonstrate accurate blood vessel reconstructions from as few as four angiogram sequences, matching the clinical workflows and outperforming state-of-the-art NeRF-based 3D reconstruction techniques. Future work includes applying our work on clinical data and reducing computation time.

### v2

@@ VI Discussion and Conclusion
We propose NeRF-CA, the first step toward a 4D Neural Radiance Field reconstruction method for X-ray coronary angiography in an extremely sparse-view setting. Our main contribution is the methodological aspects to solve the components of cardiac motion and sparse-views of the complex CA 4D reconstruction problem, given that literature has yet to propose a solution to the full clinical problem. Specifically, we propose a method that uniquely integrates scene decomposition with tailored regularization techniques to accurately decouple the CA scene in a static background and dynamic coronary artery structure. We demonstrate the effectiveness of our method both in the dynamic and sparse-view reconstruction settings. We have compared NeRF-CA to existing NeRF-based techniques for sparse-view 3D reconstruction, significantly outperforming them for CA blood vessel reconstruction. Our ablation studies have also demonstrated the need for the individual components proposed in our method. Notably, we achieve useful 4D reconstructions with four synthetic angiogram views, which are designed to capture realistic characteristics of CA data, showing potential for clinical adaptation.
@@
Despite these advantages, we also highlight a few limitations of NeRF-CA. NeRF-CA does not address all challenges in a clinical setting. Motion-related challenges that are not addressed include respiratory motion, contrast inhomogeneity, and synchronization across viewpoints. Specifically, our method assumes synchronous motion with the ECG signal across viewpoints, which cannot be assumed in clinical data due to these additional challenges. Future work could take inspiration from the various CA-based works that propose specific techniques to deal with these motion-related challenges in CA data [3]. Nevertheless, integrating these techniques into NeRF-CA is complex, as they typically rely on deformable models [50] or image registration [18], which require human interactions, such as segmentations. In contrast, NeRF-CA implicitly models motion without segmentations, making direct integration non-trivial and beyond the scope of this work. Furthermore, we validate our method using synthetic data, which has limitations such as simplified motion and homogeneous vessel contrast. However, this approach enabled a quantitative evaluation of NeRF-CA and its components. Given a further developed method which addresses all challenges in real CA data, testing on real data would be a necessary step. Additionally, geometry inaccuracies from C-arm calibration in clinical settings, unlike the perfect camera geometry assumed in our work, could impact performance and require further investigation. Finally, improving the running time of our method from hours to minutes is crucial for integration into clinical workflows. Future work could explore techniques like volumetric representations [36] or 3D Gaussian Splatting [51], which have shown promise in accelerating NeRF-based medical reconstructions. However, integrating these techniques is not straightforward, given that these methods do not represent the scene as a continuous space, which we rely on in our approach, both for static-dynamic decomposition and sparse-view modelling. Therefore, these existing frameworks would require large adaptations and are beyond the scope of this work.
@@
In this work, we proposed NeRF-CA, the first step toward a 4D reconstruction method for CA data in the extremely sparse-view setting without user interaction. We demonstrate adequate blood vessel reconstructions from as few as four angiogram viewswith synthetic data, showing the potential for clinical applications and outperforming state-of-the-art NeRF-based 3D reconstruction techniques.
Acknowledgements
This research was performed within the MEDUSA project as part of the Eindhoven MedTech Innovation Center collaboration between Eindhoven University of Technology, Philips Medical Systems, and Catharina Ziekenhuis Eindhoven.